OpenStudy (anonymous):

find tne expression and write in standard form form (square root of 24)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

\[\sqrt{24}=\sqrt{4*6}=\sqrt{4}{\sqrt6}=2\sqrt{6}\]